### v1

# Effects of tidal charge on Blandford – Znajek process around braneworld black holes

Abstract
The Blandford-Znajek (BZ) process is a pivotal mechanism to efficiently extract the energy from a rotating black hole (BH) via its plasma-filled magnetosphere in relativistic astrophysics. Within the framework of extended BZ monopole expansion, we have studied BZ process in the Randall-Sundrum braneworld BH spacetime and analyzed effects of the tidal charge on the energy and angular momentum extraction rates. It is found that the positive tidal charge reduces the BZ power of a braneworld BH, while the negative tidal charge enhances the power. Compared with a Kerr BH of the same mass and angular velocity, the BZ power exhibits a maximum reduction of approximately $15.2\%$ in positive cases, whereas in negative cases, it achieves a maximum enhancement of $66.5\%$ in power output. A similar qualitative trend is also observed for the relative angular momentum extraction rate, albeit with different magnitudes.
pacs:
04.70. Cs, 98.62.Mw, 97.60.Lf
I Introduction
When a rotating BH is immersed in a magnetic field, Blandford and Znajek demonstrated in their seminal work Blandford:1977ds that the energy and angular momentum can be electromagnetically extracted from the BH. Currently, the BZ mechanism stands as the leading theoretical model for explaining relativistic jets launched by supermassive BHs at the center of galaxies, with supporting evidence from the Event Horizon Telescope observations of M87* EventHorizonTelescope:2019pgp ; EventHorizonTelescope:2021srq .
In general, one can determine the BZ power for a given rotating BH via a regular solution of the force-free Maxwell equations in the BH spacetime. By means of perturbative techniques, the extraction rate was initially established to exhibit a quadratic scaling with the BH angular velocity, $\Omega_{\text{H}}$. However, both analytical Camilloni:2022kmx and numerical Tchekhovskoy:2009ba studies suggest that the full BZ power expression requires the inclusion of a high-spin factor, typically of the form $f(\Omega_{\text{H}})=1+a\Omega_{\text{H}}^{2}+b\Omega_{\text{H}}^{4}+\cdots$. For a rapidly rotating BH, the higher-order correction terms become relevant. So far, the formula of $f(\Omega_{\text{H}})$ for a monopolar magnetosphere in Kerr spacetime has been analytically extended to the sixth order at most Camilloni:2022kmx , while the authors in Ref. Tchekhovskoy:2009ba note the fact that high-spin factor is actually independent of the magnetospheric geometry. Recently, Camilloni et al. Camilloni:2023wyn investigated the BZ jets in the Kerr-MOG spacetime, a rotating BH solution in scalar-tensor-vector gravity, and explicitly derived its $f(\Omega_{\text{H}})$ up to orders $\mathcal{O}(\Omega_{\text{H}}^{2})$, which reveals that those expansion coefficients ($a,b,\dots$) are theory-dependent. To be specific, the high-spin factor of a Kerr-MOG BH is always smaller than that of a Kerr BH with the same mass and angular velocity. More importantly, breaking the parameter degeneracy in $\Omega_{\text{H}}$ enables the function $f(\Omega_{\text{H}})$ to serve as a new probe for testing possible deviations from general relativity (GR) in the strong-field regime Camilloni:2023wyn . This motivates us to explore the BZ power in other non-Kerr cases.
We shall focus our attention on a rotating BH in the Randall-Sundrum braneworld scenario, wherein our universe is conceived as a 3-brane embedded in a warped five-dimensional anti-de Sitter bulk Randall:1999vf . Within this framework, Aliev and Gümrükçüoğlu Aliev:2005bi solved the effective Einstein field equations on the brane, thereby obtaining a stationary and axisymmetric solution of Kerr-Newman type with the usual $Q^{2}$ term replaced by a so-called tidal charge $b$. Consequently, the sign of the tidal charge has a great impact on the global structure of black holes, the orbital dynamics of test particles, and the observational features of such systems Pun:2008ua ; Schee:2008kz ; Stuchlik:2008fy ; Abdujabbarov:2009az ; Aliev:2009cg ; Amarilla:2011fx ; Aliev:2013jqz ; Blaschke:2016uyo ; Stuchlik:2017rir ; Vagnozzi:2019apd ; Khan:2019gco ; Banerjee:2019nnj ; Nucamendi:2019qsn ; Neves:2020doc ; Hou:2021okc ; Banerjee:2021aln ; Du:2021czy ; Wei:2022jbi ; Banerjee:2022jog ; Zi:2024dpi ; Liu:2024brf ; Kumar:2025njz ; Wang:2025fuw . It is argued that a negative value of $b$ is physically more natural choice because it represents an imprint of the gravitational effects from the bulk space with a negative cosmological constant Dadhich:2000am ; Chamblin:2000ra ; Aliev:2005bi . Moreover, significant research effort has also been devoted to wave-related phenomena in braneworld BH spacetime deOliveira:2020lzp ; deOliveira:2020jha ; Dey:2020pth ; Chakraborty:2021gdf ; Mishra:2021waw ; Biswas:2021gvq ; Bohra:2023vls . In Wei:2022jbi , Wei et al. compared the influences of tidal charge on energy extraction through magnetic reconnection and BZ mechanism. Nevertheless, the equation employed to calculate BZ power is derived for Kerr BH in the context of standard GR Tchekhovskoy:2009ba , which is not a priori valid for braneworld BH, as mentioned before. The main purpose of this work is to study the BZ process rates for Randall-Sundrum braneworld BH, with explicit consideration of the dependence of the high-spin factor on the tidal charge in the energy and angular momentum fluxes.
The manuscript is organized as follows. In the next section, we briefly review the key features of the rotating braneworld BH. To evaluate the BZ rates, in Sec. III an extended BZ perturbative approach is used to develop a monopolar force-free magnetosphere around the BH. In Sec. IV, we analyze in detail the effects of tidal charge on the total energy and angular momentum fluxes with a non-trivial high-spin factor. Finally, in Sec. V we conclude.
Unless otherwise stated, we adopt geometrized units for which $c=G=1$ and Heaviside-Lorentz units in electromagnetism.
II Rotating braneworld black hole
With a stationary and axisymmetric Kerr-Schild ansatz on the brane, the rotating BH model of interest is obtained by solving the effective gravitational field equations for an empty bulk space Aliev:2005bi
$R_{\mu\nu}=-E_{\mu\nu},$ (1)
where $R_{\mu\nu}$ is the Ricci tensor, and
$E_{\mu\nu}={}^{(5)}C_{ABCD}n^{A}n^{C}e^{B}_{\mu}e^{D}_{\nu}$ (2)
is the projected “electric part” of the five-dimensional Weyl tensor. In Boyer-Lindquist coordinates, the line element of the rotating braneworld BH takes the form Aliev:2005bi
$ds^{2}=$ $-\frac{\Delta-a^{2}\sin^{2}\theta}{\rho^{2}}dt^{2}+\frac{\rho^{2}}{\Delta}dr^{2}+\rho^{2}d\theta^{2}$
$+\frac{(r^{2}+a^{2})^{2}-\Delta a^{2}\sin^{2}\theta}{\rho^{2}}\sin^{2}\theta d\varphi^{2}$
$-2a\frac{r^{2}+a^{2}-\Delta}{\rho^{2}}\sin^{2}\theta dtd\varphi,$ (3)
with
$\rho^{2}$ $\equiv r^{2}+a^{2}\cos^{2}\theta,$ (4)
$\Delta$ $\equiv r^{2}-2Mr+a^{2}+b,$ (5)
where $M$, $a$ and $b$ are BH mass, specific angular momentum and tidal charge, respectively. It should be noticed that the parameter $b$, which has the dimension of length squared, may be either positive or negative, unlike its counterpart $Q^{2}$ in Kerr-Newman metric. As expected, the solution (II) admits the Killing vectors
$(\partial_{t})^{\mu}=(1,0,0,0),\quad(\partial_{\varphi})^{\mu}=(0,0,0,1).$ (6)
Two roots of the equation $\Delta=0$ are
$r_{\pm}=M\pm\sqrt{M^{2}-a^{2}-b}.$ (7)
For a fixed $a$, it is evident that a positive/negative $b$ leads to a smaller/larger event horizon radius. In other words, a positive tidal charge reduces the gravity of the BH, whereas a negative one enhances it. This can also be verified by checking the effective Komar mass of the spacetime,
$M_{\text{eff}}=M-\frac{b}{2r}\left(1+\frac{a^{2}+r^{2}}{ar}\tan^{-1}\frac{a}{r}\right).$ (8)
Furthermore, negative tidal charges have the potential to drive a braneworld BH into an over-rotating regime, i.e., the spin $a$ exceeds the mass $M$ (see Fig. 1), which is never allowed in Einstein theory. In spite of this, a straightforward computation shows that the BH angular velocity
$\Omega_{\text{H}}=\frac{a}{2Mr_{+}-b}$ (9)
is still bounded above by $1/2M$, just as in the Kerr family.
Figure 1: The $(a,b)$ plane for braneworld BHs. A key feature is that with a negative $b$, it is possible to set $a>M$ without developing any intrinsic singularities. The region below the horizontal dashed line and to the right of the vertical dashed line corresponds to the GR exclusion zone.
The outer infinite redshift surface is located at
$r_{\text{e}}=M+\sqrt{M^{2}-a^{2}\cos^{2}\theta-b}.$ (10)
Similarly, a negative value of $b$ extends the radius of ergosphere, implying that braneworld BHs with negative tidal charges are supposed to be more energetic objects Aliev:2005bi . We emphasize, however, that for energy extraction a larger $r_{\text{e}}$ does not inherently yield a greater amount or higher efficiency, as both of which in certain scenarios depend on the size of the ergoregion. For instance, it has been shown in deOliveira:2020lzp that the difference between $r_{\text{e}}$ and $r_{+}$ decreases with the increase of b-b, thus resulting in a smaller superradiant amplification factor for massless scalar waves in non-extremal cases.
III Monopolar magnetosphere
The extraction of rotational energy via BZ mechanism is enabled by the presence of a magnetosphere around a BH. To determine the power associated with such a process, for our purpose it is necessary to solve Maxwell’s equations in the background (II). We model the magnetosphere with a monopolar topology here because of the weak dependence of high-spin factor on the specific magnetic field configuration Tchekhovskoy:2009ba .
III.1 Force-free electrodynamics
In the framework of BZ theory Blandford:1977ds , a rotating BH is surrounded by a magnetosphere filled with plasma, whose energy density is much lower than that of the electromagnetic field. As a result, the inertia of the plasma itself can be neglected and the following force-free condition approximately holds:
$F_{\mu\nu}J^{\nu}=0,$ (11)
where $F_{\mu\nu}$ is the field intensity 2-form and $J^{\nu}$ the current density of the source. With this constraint, the Maxwell equation $\nabla_{\mu}F^{\mu\nu}=-J^{\nu}$ becomes
$F_{\mu\nu}\nabla_{\sigma}F^{\sigma\nu}=0,$ (12)
or equivalently,
$F_{\mu\nu}\partial_{\sigma}\left(\sqrt{-g}F^{\sigma\nu}\right)=0.$ (13)
Another of the Maxwell equations, the Bianchi identities $dF=0$, are automatically satisfied if we choose a gauge potential $A=A_{\mu}dx^{\mu}$ such that $F=dA$. Due to the restriction from the symmetries (6), all the components $A_{\mu}$ are only functions of $r$ and $\theta$.
To describe a stationary and axisymmetric magnetosphere, in the force-free limit it is convenient to introduce three scalar functions that are invariant along the magnetic field lines Gralla:2014yja ; Camilloni:2022kmx . The first among them is the magnetic flux (or stream function) passing through a circular loop with radius $r\sin\theta$ centered on the rotational axis of the black hole,
$\Psi=A_{\varphi}.$ (14)
The second one is the angular velocity of the magnetic field lines
$\Omega=-\frac{\partial_{r}A_{t}}{\partial_{r}A_{\varphi}}=-\frac{\partial_{\theta}A_{t}}{\partial_{\theta}A_{\varphi}}.$ (15)
And the third is the poloidal current
$I=\sqrt{-g}F^{\theta r}.$ (16)
However, these quantities are not completely independent. Their definitions, together with Eq. (13), yield the integrability conditions Camilloni:2022kmx
$(\partial_{r}\Omega)(\partial_{\theta}\Psi)$ $=(\partial_{\theta}\Omega)(\partial_{r}\Psi),$ (17)
$(\partial_{r}I)(\partial_{\theta}\Psi)$ $=(\partial_{\theta}I)(\partial_{r}\Psi),$ (18)
which imply that the gradients of $\Omega$ and $I$ are everywhere parallel to the gradient of $\Psi$, reducing the functional relations to $\Omega=\Omega(\Psi)$ and $I=I(\Psi)$. In a physical picture, each magnetic field line rigidly rotates with a single angular velocity $\Omega(\Psi)$, while a fixed current $I(\Psi)$ flows along it, both being determined globally by the magnetic flux $\Psi$ the line carries.
Given the above, the general form of a force-free electromagnetic field in the spacetime (II) can be expressed as Gralla:2014yja
$F=d\Psi\wedge\eta-I\frac{\rho^{2}}{\Delta\sin\theta}dr\wedge d\theta,$ (19)
where the 1-form
$\eta=d\varphi-\Omega dt$ (20)
is co-rotating with the magnetic field lines. With the help of (19), we may further turn the equations (13) into Camilloni:2022kmx
$\eta_{\mu}\partial_{r}\left(\eta^{\mu}\Delta\sin\theta\partial_{r}\Psi\right)$ $+\eta_{\mu}\partial_{\theta}\left(\eta^{\mu}\sin\theta\partial_{\theta}\Psi\right)$
$+\frac{\rho^{2}}{\Delta\sin\theta}I\frac{dI}{d\Psi}=0,$ (21)
where $\sqrt{-g}=\rho^{2}\sin\theta$ was used. A physically acceptable solution to the above equation must be regular at the event horizon and asymptotic spacelike infinity, which necessitates the Znajek conditions Znajek:1977 ; Komissarov:2004ms ; Nathanail:2014aua ; Gralla:2014yja ; Armas:2020mio ; Camilloni:2022kmx
$I\left(r_{+},\theta\right)$ $=\left[\frac{r^{2}+a^{2}}{\rho^{2}}\sin\theta\left(\Omega_{\text{H}}-\Omega\right)\partial_{\theta}\Psi\right]\hskip-1.0pt\bigg|_{\hskip 0.0ptr_{+}},$ (22)
$I\left(\infty,\theta\right)$ $=\sin\theta\,\Omega\left(\infty,\theta\right)\partial_{\theta}\Psi\left(\infty,\theta\right).$ (23)
In addition, the field lines must smoothly across the critical surfaces defined by $\eta^{\mu}\eta_{\mu}=0$, namely
$g_{tt}+2\Omega g_{t\varphi}+\Omega^{2}g_{\varphi\varphi}=0.$ (24)
As shown in Komissarov:2004ms : (i) this equation admits only two real, distinct roots, the smaller of which is called the inner light surface (ILS) and the larger of which is called the outer light surface (OLS); (ii) the ILS is closed and always lies between the event horizon $r_{+}$ and the ergosurface $r_{\text{e}}$, whereas the OLS exhibits a cylindrical topology and does not intersect the ILS at all; (iii) in the non-rotating limit, the ILS coincides with the event horizon, and the OLS recedes to infinity.
III.2 Static background solutions
It is difficult to directly solve Maxwell’s equations in a rotating background. However, for small spin $a$, an analytic magnetosphere may be obtained by means of BZ perturbative techniques Blandford:1977ds . To this end, we first need a leading-order solution as a seed.
In the static case, $a=0$, from Eq. (II) it follows that
$ds^{2}=-\frac{\bar{\Delta}}{r^{2}}dt^{2}+\frac{r^{2}}{\bar{\Delta}}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\varphi^{2},$ (25)
where $\bar{\Delta}=(r-\bar{r}_{-})(r-\bar{r}_{+})$, $\bar{r}_{\pm}=M\left(1\pm\sqrt{1-\beta}\right)$, and $\beta\equiv b/M^{2}$. With this metric, both the angular velocity of the magnetic field lines $\Omega$ and the current $I$ vanish, the stream equation (III.1) reduces to
$\mathcal{L}\Psi(r,\theta)=0,$ (26)
where the self-adjoint operator
$\mathcal{L}=\frac{1}{\sin\theta}\frac{\partial}{\partial r}\left(\frac{\bar{\Delta}}{r^{2}}\frac{\partial}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial}{\partial\theta}\left(\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\right).$ (27)
For the split-monopole configurations that are symmetric relative to the equatorial plane, the flux must conform to the boundary conditions Nathanail:2014aua ; Armas:2020mio ; Camilloni:2022kmx
$\Psi(r,0)=1,\quad\Psi(r,\pi/2)=1,\quad\partial_{\theta}\Psi(r,0)=0.$ (28)
and must be finite as $r\to\infty$.
Assuming a solution of the form
$\Psi=c_{0}+\sum_{\ell}R_{\ell}(r)\Theta_{\ell}(\theta),$ (29)
Eq. (26) can be decoupled into two ordinary differential equations:
$\frac{d}{dr}\left(\frac{\bar{\Delta}}{r^{2}}\frac{dR_{\ell}}{dr}\right)-\frac{\ell(\ell+1)}{r^{2}}R_{\ell}$ $=0,$ (30)
$\frac{d}{d\theta}\left(\frac{1}{\sin\theta}\frac{d\Theta_{\ell}}{d\theta}\right)+\frac{\ell(\ell+1)}{\sin\theta}\Theta_{\ell}$ $=0.$ (31)
Obviously, the tidal charge $\beta$ remains only in the radial part, while the regular solution to the angular equation (31) is the same as that in the Schwarzschild spacetime, which is given by Gralla:2015vta
$\Theta_{\ell}=\begin{cases}{}_{2}F_{1}\left(\dfrac{\ell}{2},-\dfrac{\ell+1}{2},\dfrac{1}{2};\cos^{2}\theta\right),&\ell\text{ odd}\vskip 5.0pt\\ {}_{2}F_{1}\left(-\dfrac{\ell}{2},\dfrac{\ell+1}{2},\dfrac{3}{2};\cos^{2}\theta\right)\cos\theta,&\ell\text{ even}\end{cases},$ (32)
where $_{2}F_{1}$ denotes the hypergeometric function. For the lowest order ones, e.g.,
$\Theta_{0}=\cos\theta,\quad\Theta_{1}=\sin^{2}\theta,$ (33)
$\Theta_{2}=\cos\theta\sin^{2}\theta,\quad\Theta_{3}=\left(1-5\cos^{2}\theta\right)\sin^{2}\theta.$ (34)
On the other hand, Camilloni et al. Camilloni:2023wyn reformulated Eq. (30) as a Heun’s equation characterized by four regular singular points and showed that its general solution consists in a linear combination of two sets of eigenfunctions,
$R_{\ell}=c_{1}U_{\ell}(w_{\beta};w)+c_{2}V_{\ell}(w_{\beta};w),\quad\ell\geqslant 1,$ (35)
where the dimensionless variables
$w=\frac{r}{\bar{r}_{+}},\quad w_{\beta}=\frac{\bar{r}_{-}}{\bar{r}_{+}}.$ (36)
The explicit expressions for $U_{\ell}$ and $V_{\ell}$ can be easily found in Ref. Camilloni:2023wyn , and their asymptotic behaviors at the static event horizon and at spatial infinity read
$U_{\ell}(w_{\beta};1)$ $\sim(2M)^{\ell+1}(1-w_{\beta})^{\ell},$ (37)
$V_{\ell}(w_{\beta};1)$ $\sim(2M)^{-\ell}\log\left(\frac{w-1}{w-w_{\beta}}\right),$ (38)
$U_{\ell}(w_{\beta};\infty)$ $\sim(2Mw)^{\ell+1},$ (39)
$V_{\ell}(w_{\beta};\infty)$ $\sim(2Mw)^{-\ell}.$ (40)
The radial harmonics (35) are valid for any Reissner-Nordström type metric like (25). Collecting these results above, it is clear that the homogeneous equation (26) does not have a general solution that is regular over the entire space.
III.3 Perturbed configurations
The extended BZ method allows the force-free field variables to be expanded as a power series in the BH spin parameter $\alpha\equiv a/M$ Armas:2020mio ; Camilloni:2022kmx :
$\Psi$ $=\sum_{n=0}^{\infty}\alpha^{n}\Psi_{n}(r,\theta),$ (41)
$\Omega$ $=\frac{1}{2M}\sum_{n=0}^{\infty}\alpha^{n}\Omega_{n}(r,\theta),$ (42)
$I$ $=\frac{1}{2M}\sum_{n=0}^{\infty}\alpha^{n}I_{n}(r,\theta).$ (43)
The factor $2M$ is included so that $\Omega_{n}$ and $I_{n}$ are dimensionless. Henceforth, we follow Ref. Camilloni:2023wyn and retain terms up to order $\alpha^{3}$.
As a seed solution, our starting point is the split-monopole configuration
$\Psi_{0}=1-\cos\theta,\quad\Omega_{0}=I_{0}=0,$ (44)
where the polar angle ranges between $0,\pi/2]$. It is easy to verify that $\Psi_{0}$, which is consistent with the boundary conditions (28), is indeed a vacuum solution to Eq. (26).
At the first order in $\alpha$, the integrability conditions (17) and (18) indicate that neither $\Omega_{1}$ nor $I_{1}$ is a function of $r$, meaning
$\Omega_{1}=\omega_{1}(\theta),\quad I_{1}=i_{1}(\theta).$ (45)
Substituting Eqs. (41), (42) and (43) into (III.1) and keeping only up to the first order terms in the perturbations, we find
$\mathcal{L}\Psi_{1}(r,\theta)=0,$ (46)
where the operator $\mathcal{L}$ is defined by Eq. (27). By virtue of the asymptotic behaviors (37)–(40), as was mentioned previously, regularity at both the horizon and infinity forces the solution to be trivial
$\Psi_{1}=0.$ (47)
Then, the Znajek conditions (22) and (23) at the first order in $\alpha$ give
$\omega_{1}=\frac{1}{\left(1+\sqrt{1-\beta}\right)^{2}},\quad i_{1}=\omega_{1}\Theta_{1}.$ (48)
Moving to the next perturbative order, it turns out that $\Omega_{2}=\omega_{2}(\theta)$, $I_{2}=i_{2}(\theta)$, and $\Psi_{2}$ obeys the sourced equation
$\mathcal{L}\Psi_{2}=-\frac{4M^{3}}{r^{4}\bar{r}_{+}^{2}}\left(r-\frac{\bar{r}_{+}\bar{r}_{-}}{2M}\right)\frac{r+\bar{r}_{+}}{r-\bar{r}_{-}}\frac{\Theta_{2}}{\sin\theta}.$ (49)
A particular solution to this equation is found to be $\Psi_{2}=R_{2}(r)\Theta_{2}(\theta)$ with Camilloni:2023wyn
[…]
$-\frac{w\left(1+w_{\beta}\right)\left(3+w_{\beta}\right)}{2}-\frac{w_{\beta}\left[17+w_{\beta}\left(20-w_{\beta}\right)\right]}{9}$
$-\frac{6w\left[4w-\left(1+w_{\beta}\right)\right]-\left[1+w_{\beta}\left(10+w_{\beta}\right)\right]}{6}$
$\times\left[\log\left(w\right)-\left(1+w_{\beta}\right)\log\left(\frac{w-w_{\beta}}{1-w_{\beta}}\right)\right]$
[…]
$-\operatorname{Li}_{2}\left(\frac{1}{w}\right)+\log\left(w\right)\log\left(\frac{w-1}{w-w_{\beta}}\right)$
$+\left(1+w_{\beta}\right)\left(\frac{\pi^{2}}{6}+\frac{1}{2}\log^{2}\left(\frac{w-w_{\beta}}{1-w_{\beta}}\right)-\frac{1-w_{\beta}}{2w_{\beta}}\right$.
$\times\left.\left.\left.\!\!\log\left(1-\frac{w_{\beta}}{w}\right)+\operatorname{Li}_{2}\left(\frac{1-w}{1-w_{\beta}}\right)\right)\right]\right\},$ (50)
where $\operatorname{Li}_{2}$ denotes the dilogarithm function. Recall that parameter $w_{\beta}$ is linked to the tidal charge via $w_{\beta}=(1-\sqrt{1-\beta})/(1+\sqrt{1-\beta})$, in the limit of $w_{\beta}\to 0$, Eq. (50) reduces to its corresponding form derived for the Kerr geometry Blandford:1977ds ; Camilloni:2022kmx ,
$R_{2}={}$ $\frac{11}{72}+\frac{1}{6w}+w\left(1-2w\right)+\frac{1+6w-24w^{2}}{12}$
$\times\log\left(w\right)+\frac{w^{2}\left(4w-3\right)}{2}\left[\operatorname{Li}_{2}\left(\frac{1}{w}\right)\right$.
$\left.-\log\left(w\right)\log\left(1-\frac{1}{w}\right)\right],$ (51)
as a check. These kinds of solutions are obtained by using the Green’s function method. For details, see, e.g., Appendix B in Ref. Camilloni:2022kmx . To visualize the behavior of the function $R_{2}(w)$ we plot it in Fig. 2, which also compares it with the standard GR case (51) marked by a black solid line.
Figure 2: $R_{2}$ versus the dimensionless radial coordinate $w$ for different values of $\beta$. All of the curves seamlessly traverse the static event horizon $w=1$ and run towards zero as $w\to\infty$.
Further, the Znajek conditions (22) and (23) tell us that
$\omega_{2}=i_{2}=0.$ (52)
At the order of $\alpha^{3}$, the integrability conditions (17) and (18) lead to $\Omega_{3}=\omega_{3}(\theta)$, and
$I_{3}=\frac{\partial_{\theta}i_{1}}{\partial_{\theta}\Psi_{0}}\Psi_{2}(r,\theta)+i_{3}(\theta).$ (53)
From the stream equation (III.1) we get $\mathcal{L}\Psi_{3}(r,\theta)=0$, which implies, once again, that
$\Psi_{3}=0.$ (54)
Finally, the Znajek conditions (22) and (23) produce
$\omega_{3}={}$ $\frac{1}{\left(1+\sqrt{1-\beta}\right)^{4}}\left\{\frac{1}{\sqrt{1-\beta}}+\frac{1}{2}\right$.
$\times\left.\left[1-\left(1+\sqrt{1-\beta}\right)^{2}R_{2}^{\text{H}}\right]\Theta_{1}\right\},$ (55)
[…]
$+\left.\frac{1}{10}\left[\frac{1}{\left(1+\sqrt{1-\beta}\right)^{2}}-R_{2}^{\text{H}}\right]\Theta_{3}\right\},$ (56)
where $R_{2}^{\text{H}}$ is defined as the function (50) evaluated at the static event horizon $w=1$.
With several $\Omega_{n}$ in hand, we may now determine the light-surface positions perturbatively by solving the algebraic equation (24). For ILS, the result is
$\frac{r_{\text{ILS}}}{M}=x_{0}+\alpha^{2}x_{2}+\alpha^{4}x_{4}+\mathcal{O}\left(\alpha^{5}\right),$ (57)
in which the coefficients are
$x_{0}$ $=1+\sqrt{1-\beta},$ (58)
$x_{2}$ $=-\frac{1}{2\sqrt{1-\beta}}\left(1-\frac{\Theta_{1}}{4}\right),$ (59)
[…]
$-\left[1+\frac{8}{1+\sqrt{1-\beta}}+16\left(1-\beta\right)R_{2}^{\text{H}}\right]\frac{\Theta_{1}}{20}$
$\left.+\left[9-\frac{8}{1+\sqrt{1-\beta}}-16\left(1-\beta\right)R_{2}^{\text{H}}\right]\frac{\Theta_{3}}{80}\right\}.$ (60)
Similarly for the outer one,
$\frac{r_{\text{OLS}}}{M}=\frac{y_{-1}}{\alpha}+y_{0}+\alpha y_{1}+\mathcal{O}\left(\alpha^{2}\right),$ (61)
[…]
The relations between the two light surfaces and the black hole spin are illustrated in Fig. 3. For a fixed tidal charge $\beta$, the values of both $r_{\text{ILS}}$ and $r_{\text{OLS}}$ decrease as $\alpha$ increases. What differs, however, is that $r_{\text{ILS}}$ begins its drop at the static horizon $\bar{r}_{+}$, whereas $r_{\text{OLS}}$ is dragged in from infinity to a finite distance by the rotation. For a given $\alpha$, a positive $\beta$ leads to smaller $r_{\text{ILS}}$ and $r_{\text{OLS}}$, while a negative $\beta$ has the opposite effect.
Figure 3: Plots of $r_{\text{ILS}}$ and $r_{\text{OLS}}$ as functions of the dimensionless spin parameter $\alpha$ for selected values of $\beta$, with $\theta=30^{\circ}$. The right ends of the lines are cut off because for each $\beta$, there exists a physical bound $\sqrt{1-\beta}$ on $\alpha$, as shown in Fig. 1. In the left panel, the gray shading delineates the extent of the ergoregion at the same perturbative order as Eq. (57).
Fig. 4 shows the obtained monopolar magnetospheres around braneworld BHs in different rotating regimes — specifically, $\alpha<1$, $\alpha=1$, and $\alpha>1$. Observe that all the magnetic field lines escape from the event horizon, extend to spatial infinity, and cross the critical surfaces in a well-behaved manner. Compared with the initial configuration, the high-order field lines have a tendency to bunch up towards the BH’s rotation axis. The larger $\alpha$ or $\beta$ is, the closer they are to the axis, thereby enhancing the radial magnetic fields on the event horizon (see Fig. 5). In Ref. Gralla:2015uta , the bunching of field lines in BH magnetospheres with various geometries was explored analytically.
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Figure 4: Magnetic field lines as the contours of flux $\Psi$ on the $(r,\theta)$ plane for $\alpha=0.7$ (top row), $\alpha=1.0$ (middle row), and $\alpha=1.3$ (bottom row). The leading-order solution (44) is also included and depicted with dotted lines. In addition, the gray, red, and purple curves are used to identify the ergosurface, ILS, and OLS, respectively.
Figure 5: The radial magnetic field $B^{r}=\left(\partial_{\theta}\Psi\right)/\sqrt{-g}$ at the north pole of the braneworld BH horizon. A dashed line defines the boundary of physically acceptable configurations.
IV Blandford – Znajek rates
For a stationary and axisymmetric magnetosphere, there are two Noether currents associated with the Killing fields (6), i.e.,
$\mathcal{J}^{E}_{\mu}=-T_{\mu\nu}(\partial_{t})^{\nu},\quad\mathcal{J}^{L}_{\mu}=T_{\mu\nu}(\partial_{\varphi})^{\nu},$ (62)
where the stress-energy tensor is of course
$T_{\mu\nu}=F_{\mu\sigma}F_{\nu}{}^{\sigma}-\frac{1}{4}g_{\mu\nu}F_{\sigma\rho}F^{\sigma\rho}.$ (63)
Physically, $\mathcal{J}^{E}_{\mu}$ is the energy flux density and $\mathcal{J}^{L}_{\mu}$ the angular momentum flux density. Integrating the dual form of $\mathcal{J}^{E}_{\mu}$ over a hypersurface generated by a sphere $S^{2}$ of radius $r$ cross a lapse $\Delta t$ of time, in the $\Delta t\to 0$ limit, gives the total flux of energy flowing out of that sphere Gralla:2014yja ; Camilloni:2022kmx ,
$\dot{E}=2\pi\int_{0}^{\pi}\Omega(r,\theta)I(r,\theta)\partial_{\theta}\Psi(r,\theta)d\theta,$ (64)
where the overdot stands for the derivatives with respect to the Killing time $t$. Similarly, the outward angular momentum flux reads Gralla:2014yja ; Camilloni:2022kmx
$\dot{L}=2\pi\int_{0}^{\pi}I(r,\theta)\partial_{\theta}\Psi(r,\theta)d\theta.$ (65)
The energy and angular momentum extraction rates from a rotating BH via the BZ process are determined by Eqs. (64) and (65), respectively. Since the fluxes are conserved, one can evaluate them at the event horizon $r=r_{+}$.
Feeding the invariants (41)–(43) into Eq. (64), the power is found to be
$\dot{E}=\alpha^{2}\dot{E}_{2}+\alpha^{4}\dot{E}_{4}+\mathcal{O}\left(\alpha^{6}\right),$ (66)
with
$\dot{E}_{2}={}$ $\frac{2\pi}{3M^{2}\left(1+\sqrt{1-\beta}\right)^{4}},$ (67)
$\dot{E}_{4}={}$ $\frac{4\pi}{15M^{2}\left(1+\sqrt{1-\beta}\right)^{6}}\left\{\frac{5}{\sqrt{1-\beta}}\right$.
+ $\left.\ \,+\,2\left[1-\left(1+\sqrt{1-\beta}\right)^{2}\!R_{2}^{\text{H}}\right]\right\}.$ (68)
Along the same lines, from Eq. (65) we obtain
$\dot{L}$ $=\alpha\dot{L}_{1}+\alpha^{3}\dot{L}_{3}+\mathcal{O}\left(\alpha^{5}\right),$ (69)
$\dot{L}_{1}$ $=2M\left(1+\sqrt{1-\beta}\right)^{2}\dot{E}_{2},$ (70)
$\dot{L}_{3}$ $=M\left(1+\sqrt{1-\beta}\right)^{2}\dot{E}_{4}.$ (71)
If we set $\beta=0$, Eqs. (66) and (69) reduce to the energy and angular momentum fluxes derived in Kerr spacetime Tanabe:2008wm ; Camilloni:2022kmx , at their respective orders.
To find out the high-spin factor, with the help of Eq. (9) we can recast the power (66) as an expansion in the BH angular frequency. In fact, it has been argued in Tchekhovskoy:2009ba that an $\Omega_{\text{H}}$-expansion for the BZ rates is much more natural and accurate than the $\alpha$-expansion was before. Up to the fourth order in $\Omega_{\text{H}}$, we arrive at
$\dot{E}=\frac{2}{3}\pi\Omega_{\text{H}}^{2}f_{\beta}\left(\Omega_{\text{H}}\right),$ (72)
where
$f_{\beta}=1$ $+\frac{4}{5}M^{2}\Omega_{\text{H}}^{2}\left(1+\sqrt{1-\beta}\right)^{2}$
$\times\left[1-\left(1+\sqrt{1-\beta}\right)^{2}\!R_{2}^{\text{H}}\right]+\mathcal{O}\left(\Omega_{\text{H}}^{4}\right).$ (73)
The subscript $\beta$ serves to highlight the background dependence of the high-spin factor, as noted in Camilloni:2023wyn . In the $\beta\to 0$ limit,
$f_{0}=1+\frac{8}{45}\left(67-6\pi^{2}\right)M^{2}\Omega_{\text{H}}^{2}+\mathcal{O}\left(\Omega_{\text{H}}^{4}\right),$ (74)
the BZ4 formula given in Ref. Tchekhovskoy:2009ba is recovered by aligning the relevant coefficients carefully. For the angular momentum flux, it follows from Eq. (69) that
$\dot{L}=\frac{2}{3}\pi\Omega_{\text{H}}\left(1+f_{\beta}\right).$ (75)
To the best of our knowledge, Eqs. (72) and (75) are also valid for other neutral, single-parameter modified Kerr BHs with a monopolar magnetosphere. Moreover, in the absence of the high-spin factor, the remaining $\dot{E}=2\pi\Omega_{\text{H}}^{2}/3$ and $\dot{L}=4\pi\Omega_{\text{H}}/3$ are confirmed to be exactly the same as the leading order results obtained in standard GR Camilloni:2022kmx . That is to say, the BZ mechanism cannot distinguish a Kerr BH from a braneworld one under consideration unless the high-spin factor (IV) is measured, due to the degeneracy of the spin parameter $a$ and the tidal charge $b$ in Eq. (9). Indeed, the degeneracy in the leading-order term of the BZ power has been confirmed in a general stationary and axisymmetric BH spacetime Konoplya:2021qll .
What is really of concern, as suggested in Ref. Camilloni:2023wyn , is the fractional deviation of $f_{\beta}$ with respect to its GR limit $f_{0}$, or equivalently, is the relative difference in BZ power between a Kerr and a braneworld BH with the same mass $M$ and angular velocity $\Omega_{\text{H}}$,
$\Delta\dot{E}=\frac{f_{\beta}-f_{0}}{f_{0}}.$ (76)
It is clear from Fig. 6 that a positive $\beta$ always leads to a negative $\Delta\dot{E}$, and therefore tends to reduce the relative energy extraction rate. With a high-spin factor accurate to the second order in $\Omega_{\text{H}}$, the effect of positive tidal charges can suppress the BZ power of a braneworld BH by up to almost $15.2\%$ compared to that of a Kerr BH. Conversely, a negative tidal charge acts to boost the power by $66.5\%$ at most from a purely theoretical perspective. Such an enhancement, however, can only be achieved at $\beta\approx-58.8$, a value that is likely not allowed in realistic astrophysical contexts. For example, observations of the BH shadow constrain the dimensionless tidal charge to $\beta=-0.1^{+0.6}_{-0.5}$ Banerjee:2019nnj ; Banerjee:2022jog ; Krishnendu:2024jkj , corresponding to a maximum $\Delta\dot{E}$ of $11.8\%$.
Figure 6: Percentage-scaled $\Delta\dot{E}$ as a function of the dimensionless angular velocity $\omega_{\text{H}}\equiv M\Omega_{\text{H}}$ for different values of $\beta$, where the independent variable is bounded by $\sqrt{1-\beta}/(2-\beta)$ according to Eq. (9). Since there are no allowed configurations outside the dashed gray line, $\Delta\dot{E}$ reaches its minimum of approximately $15.2\%$ at $\omega_{\text{H}}=0.392$ and $\beta=0.765$. The inset shows the envelope at extremely negative tidal charges, revealing a maximum enhancement $\Delta\dot{E}_{\text{max}}\approx 66.5\%$ for $\omega_{\text{H}}=0.127$ with $\beta=-58.818$.
And then we finally turn to the relative angular momentum flux,
$\Delta\dot{L}=\frac{f_{0}}{1+f_{0}}\Delta\dot{E},$ (77)
which behaves qualitatively similarly to Eq. (76), as presented in Fig. 7.
Figure 7: Relative difference in angular momentum extraction rate, $\Delta\dot{L}$, between Kerr and braneworld Kerr BHs with identical $M$ and $\Omega_{\text{H}}$. The function (77) exhibits a maximum of $33.7\%$ at $\omega_{\text{H}}=0.143$, $\beta=-45.602$, and a minimum of $8.3\%$ at $\omega_{\text{H}}=0.401$, $\beta=0.749$.
V Conclusions
This work provides an analytical investigation into the effects of tidal charge on the BZ process rates in a rotating braneworld BH spacetime. To do so, following Ref. Camilloni:2023wyn , we expand the variables $\Psi$, $\Omega$, and $I$ as a truncated power series in the dimensionless spin parameter $\alpha$, and thereby obtain a monopolar BH magnetosphere that is regular at all the critical surfaces by solving the force-free Maxwell equations perturbatively. For higher precision, we recast the BZ rates as expansions in the BH angular velocity $\Omega_{\text{H}}$, which yields a high-spin factor manifestly dependent on the tidal charge.
For Kerr and braneworld BHs sharing the same mass $M$ and angular velocity $\Omega_{\text{H}}$, it is found that the BZ power is suppressed by the positive tidal charge, with a maximum reduction of approximately $15.2\%$, whereas a negative tidal charge can enhance the power by $66.5\%$ at most theoretically. A similar trend also holds for the angular momentum extraction rate, which is suppressed by a positive tidal charge and enhanced by a negative one, with the extremal values of $8.3\%$ and $33.7\%$, respectively. These results are helpful in placing independent constraints on the tidal charge parameter and the extra dimension size through future horizon-scale observations.
A direct extension of our work would be to incorporate higher-order contributions to the high-spin factor for the braneworld BH, alongside a comparative study of its BZ process with other known energy extraction mechanisms.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant No.12275078, 11875026, 12035005, 2020YFC2201400 and the innovative research group of Hunan Province under Grant No. 2024JJ1006.